A triangular windowpane is 4 feet 8 inches wide and 3 feet 6 inches high. What is the area of the windowpane? If necessary, round to the nearest tenth.
Answer by fcabanski(1391)

The area of a triangle is 1/2 * b * h where b is the width of the base, and h is the height.


4 feet 8 inches is 4 and 2/3 feet or 14/3 feet and 3 feet 6 inches is 3 and 1/2 feet or 7/2 feet.


1/2 * 14/3 * 7/2 = 98/12 = 8 and 2/12 feet or 8 1/6 feet or 8 feet 2 inches.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)